Sheridan Company issues $6070000 face value of bonds at 95 on January 1, 2016. The bonds are dated January 1, 2016, pay interest semiannually at 8% on June 30 and December 31, and mature in 10 years. Straight-line amortization is used for discounts and premiums. On September 1, 2019, $3642000 of the bonds are called at 101 plus accrued interest. What gain or loss would be recognized on the called bonds on September 1, 2019?

Entry field with incorrect answer

$260050 loss.
$364200 loss.
$151750 loss.
$204550 loss.
Answer #1

Discount on Bonds = 6070000*(1-0.95) = $303500

Amortization of discount upto September 1, 2019 = 303500*44/120 = 111283

Balance in discount account = 303500 - 111283 = $192217

Discount related to called bonds = 192217*3642000/6070000 = $115330

Loss recognized on called bonds = discount + premium at the time of payment

= $115330+3642000*1%

= $151750 Loss